# A Survey on RAG Meets LLMs: Towards Retrieval-Augmented Large Language Models

### v1

@@ 3.1.3. Pre-retrieval and Post-retrieval Enhancement
To ensure the retrieval quality, i.e., increase the accuracy and relevance of the retrieved results, various pre- and post-retrieval strategies have been proposed to further enhance the input and output of the retriever. Wang et al. (2023f) propose a query expansion approach Query2doc, which generates pseudo-documents by few-shot prompting LLMs and expands the query with the relevant information in pseudo-documents, which can aid in query disambiguation and guide the retrievers. They have empirically demonstrated that such a method can boost the performance of both the sparse and dense retriever (Karpukhin et al., 2020) on ad-hoc information retrieval datasets. Similarly, Gao et al. (2023a) propose Hypothetical Document Embedding (HyDE) method, which instructs an LLM to generate hypothetical documents for the given query. The hypothetical documents are then used as new queries to get embedded and search for neighbors with the dense retriever.
Another pre-retrieval strategy, query rewrite (Ma et al., 2023a), aims to close the gaps between the input text and the needed knowledge in retrieval, to reformulate the original question into a more conducive version to retrieve. Specifically, Ma et al. (2023a) propose the Rewrite-Retrieve-Read framework, which prompts an LLM to generate the query for the retrieval function. The motivation of the rewriting step is to clarify the retrieval need in the new query to ease the burden on the retrieval function to comprehend the input and enhance the output, i.e., retrieved relevant information. They have tested both the settings of using a frozen LLM and a trainable model to be the rewriter, both outperforming naive RAG or generation models, demonstrating diverse performance on different tested QA datasets though.
@@
More recently, advanced RAG pipelines have been proposed using LLMs to generate reasoning paths and plans with the Information Retrieval (IR) module to iteratively retrieve knowledge to enhance LLM-based generation (Yao et al., 2023; Xu et al., 2023a; Shao et al., 2023). However, Zhu et al. (2023) point out that if the outputs of IR and LLM are low-quality, the retrieval and generation processes will get hindered by each other with such an iterative guidance pipeline. To overcome this barrier, they propose a new reasoning approach for query and retrieved knowledge enhancement. Post-retrieval strategies may also function to enhance the compatibility between the retrieved results and the generation models. For example, one of the main limitations of existing LLMs is the length of the input tokens, which prevents long retrieved documents being directly incorporated into existing RA-LLMs. For this limitation, Xu et al. (2023b) propose Retrieve, Compress, Prepend (RECOMP), which adds an intermediate step to process the retrieved documents into a textual summary before in-context augmentation in the generation process.
@@ 3.1.4. Database
Retrieval in RAG is conducted based on external knowledge source, which can be a closed- or open-sourced (Ma et al., 2023a; Menick et al., 2022), as illustrated in Figure 2. Closed-sourced database generally stores key-value pairs for knowledge, which can be constructed in various ways. The keys are primarily used for similarity matching, being as sparse vectors such as in BM25 or dense embeddings from the retrieval encoding. The value depends on the specific retrieval target, which is raw text in most cases (Guu et al., 2020; Lewis et al., 2020c; Izacard and Grave, 2021b; Borgeaud et al., 2022; Lewis et al., 2020a; Seo et al., 2019). For example, each Wikipedia article is split into disjoint 100-word chunks, to make a total of 21M documents in early RAG (Lewis et al., 2020c). Each document is encoded by a dense embedding and saved in the database as the value and key, respectively. The value can store tokens too, one for each as applied in kNN-LM and Spalm (Yogatama et al., 2021; Khandelwal et al., 2020). The source of the database depends on the specific application domains and tasks. Wikipedia is one of the most commonly applied general retrieval sets in previous RAG work, which stores factual structured information and has several versions differing in scale, from billion token-level (Khandelwal et al., 2020; Yogatama et al., 2021; Lewis et al., 2020c; Guu et al., 2020; Févry et al., 2020; de Jong et al., 2022; Xu et al., 2023b; Shi et al., 2023; Ram et al., 2023) to trillion token-level (Borgeaud et al., 2022). Domain-specific database is also used for downstream tasks. For example, for the code generation task, Zan et al. (2022) collect API information and code files of public libraries to build their APIretriever database. In addition, Zhou et al (Zhou et al., 2022) propose to use a documentation pool frequently updated with new content (newly released libraries) in their model.
@@
Applying Internet searching engine (Luo et al., 2023a) such as Bing and Google avoids the maintenance of the search index and can access up-to-date knowledge (Lazaridou et al., 2022). Meanwhile, it provides a broader knowledge base than the closed-sourced database (Asai et al., 2023b; Lazaridou et al., 2022). Internet search has been widely incorporated with black-box LLMs and shows effectiveness for different functions such as knowledge augmentation (Lazaridou et al., 2022), fact-checking (Menick et al., 2022) and LLM agent enhancement (Yao et al., 2023). Compared to traditional RAG, Internet search has been leveraged more as the retriever in RA-LLMs owing to the extraordinary capability of LLMs to be the Reader to comprehend the searching results, i.e., the retrieved documents, as well as LLMs’ ability to use tools to process and analyze the them (Ma et al., 2023a). Existing studies (Yu et al., 2023a) have shown that leveraging search engines (e.g., InstrucGPT) is particularly effective for LLMs on zero-shot knowledge-intensive tasks such as OpenQA and fact checking.
@@ 3.4. Retrieval Augmentation Necessity and Frequency
Most existing methods determine the necessity of retrieval based on the preliminary answers of LLMs or their internal reasoning results (Ram et al., 2023; Min et al., 2022). For example, Self-RAG (Asai et al., 2023b) introduces special tokens to assess the necessity of retrieval and control retrieval behavior. Other methods design iterative prompts to decide if extra information is required during generation, which thereby needs to invoke retrieval or other actions for LLMs (Yao et al., 2023; Wei et al., 2022). In traditional RAGs, retrieval necessity judgment has also been explored and proposed to address by intuitive approaches such as assessing the confidence of the logits produced by the generation models (Jiang et al., 2021; Kadavath et al., 2022; He et al., 2021b). Such a solution is also applicable to RA-LLMs, for example, FLARE (Jiang et al., 2023) dynamically triggers RAG if logits are lower than a specific threshold. More flexibly, Tan et al. (2024) introduces SlimPLM, a collaborative approach that detects missing knowledge in LLMs with a slim proxy model, which functions to generate a “heuristic answer”. The “heuristic answer” is used to assess the necessity of retrieval and facilitate the retrieval process if necessary by being applied in query rewriting.
@@ 4.1. Training-free
With the huge number of parameters, LLMs have exhibited human-level intelligence and achieved promising prediction performance on various downstream tasks. However, it is extremely challenging to frequently perform fine-tuning and update the knowledge stored in the model parameters (Lewis et al., 2020c) due to the considerable time and computational resources required. Recently, numerous studies have suggested enhancing LLMs with retrieval mechanisms, enabling them to dynamically acquire new knowledge from external sources without extra training processes (i.e., training-free(Izacard and Grave, 2021b; Jiang et al., 2023; Khattab et al., 2022), instead of relying solely on the implicit knowledge encoded in the model’s parameters. These approaches have shown significant performance improvement for various knowledge-intensive tasks, such as open-domain question answering (Lewis et al., 2020c) and document summarization (Song et al., 2023). According to the different ways in which LLMs utilize retrieved information, we categorize these training-free methods into two categories: 1) Prompt Engineering-based Methods integrate retrieved knowledge into the original prompt directly, and 2) Retrieval-Guided Token Generation Methods retrieve information to calibrate the token generation process.
@@ 4.1.1. Prompt Engineering-based Methods
As the LLMs’ generation performance highly depends on the input query, numerous training-free RAG approaches employ external knowledge by refining the original prompts (Jiang et al., 2023; Khattab et al., 2022; Li et al., 2023c). Specifically, the retrieved texts are usually used as contextual information and combined with the original prompt to guide the generation of LLMs (Izacard and Grave, 2021b; Jiang et al., 2023; Khattab et al., 2022; Purwar and Sundar, 2023; Li et al., 2023c; Wang et al., 2023g; Kim et al., 2023). For example, In-Context RALM (Ram et al., 2023) keeps the LLM parameters unchanged and directly incorporates the retrieved document before the original prompt to augment the generation process. IRCoT (Trivedi et al., 2023) interleaves chain-of-thought (CoT) generation and knowledge retrieval steps, enabling the retrieval of more relevant information for subsequent reasoning steps compared to standard retrieval methods that rely solely on the question as the query. Instead of retrieving knowledge from a large corpus, GENREAD (Yu et al., 2023a) first prompts a LLM to generate contextual documents based on the query, and then generate answers based on the given context and question. SKR (Wang et al., 2023a) proposes guiding LLMs to determine whether they can answer a given question based on their internal knowledge, enabling flexible utilization of both internal and external knowledge by selectively calling the retriever. TOC (Kim et al., 2023) first retrieves relevant knowledge for ambiguous questions and recursively constructs a tree structure by clarifying ambiguous questions into multiple disambiguate questions, which is further aggregated to generate long-form answers.
@@ 5.1.1. QA Systems
QA Systems aim to provide precise answers to user’s queries. However, even when trained on extensive data, these systems may lack the latest information or specific domain knowledge that is not included in their training data (Izacard and Grave, 2021b; Liu et al., 2022b). To address this limitation, the integration of RA-LLMs has played a crucial role in advancing the capabilities of QA systems by enhancing their ability to retrieve and synthesize relevant information (Borgeaud et al., 2022; Izacard and Grave, 2021b). Specifically, RA-LLMs can provide coherent and contextually relevant answers by leveraging their retrieval component to access a vast knowledge base. For example, REALM (Guu et al., 2020) integrates a knowledge retriever that can retrieve information from a large corpus during pre-training, fine-tuning, and inference. This approach allows REALM to effectively retrieve from a vast knowledge corpus, thereby improving the accuracy of its responses. Similarly, Fusion-in-Decoder (Izacard and Grave, 2021b) retrieves passages from support documents and then fuses them with questions to generate the answer, achieving higher accuracy. In addition,  Borgeaud et al. (2022) indicate that the quality of the answers may rely more on the output of the retrieval encoder.
@@ 6. Future Challenges and Opportunities
Multi-Lingual RA-LLMs. The ability of leveraging knowledge from multiple languages can greatly enhance the capabilities of retrieval-augmented language models. As the world becomes increasingly interconnected, there is a growing need for AI systems that can understand and communicate across different languages. By incorporating multilingual knowledge retrieval and generation, these models can access and synthesize information from diverse linguistic sources, enabling more comprehensive and nuanced understanding and generation capabilities. Additionally, multilingual models can facilitate cross-cultural communication and knowledge sharing and breaking down language barriers, thereby bringing convenience to people across different regions of the world, especially those in areas with minority languages (Kabra et al., 2023; Li et al., 2023c). For example, users from countries with less prevalent languages can utilize abundant English and Chinese corpora for knowledge retrieval, enhancing the performance of large language models in downstream tasks.
@@
Quality of External Knowledge. As a commonly used datastore in current RAG systems, Wikipedia (Zhu et al., 2024; Karpukhin et al., 2020) serves as a vast repository of external textual knowledge used to augment the generation process, which contains millions of articles covering various disciplines. However, the reliability and accuracy of individual articles within Wikipedia vary significantly, and the introduction of some texts that deviate from facts might even mislead the model’s generation process. Therefore, it is crucial to enhance the quality of the external knowledge corpus and mitigate the negative impact of low-quality knowledge on the performance of LLMs. By enhancing the quality of the external knowledge and tailing robust mechanisms by filtering out low-quality or unreliable information, the RAG-empowered LLM systems might produce more accurate, reliable outputs, thereby improving their effectiveness in various real-world applications.

### v2

@@ 3.1.3. Pre-retrieval and Post-retrieval Enhancement
To ensure the retrieval quality, i.e., increase the accuracy and relevance of the retrieved results, various pre- and post-retrieval strategies have been proposed to further enhance the input and output of the retriever. Wang et al. (2023f) propose a query expansion approach Query2doc, which generates pseudo-documents by few-shot prompting LLMs and expands the query with the relevant information in pseudo-documents to improve the query disambiguation and guide the retrievers. They have empirically demonstrated that such a method can boost the performance of both the sparse and dense retriever (Karpukhin et al., 2020) on ad-hoc information retrieval datasets. Similarly, Gao et al. (2023a) propose Hypothetical Document Embedding (HyDE) method, which instructs an LLM to generate hypothetical documents for the given query. The hypothetical documents are then used as new queries to get embedded and search for neighbors with the dense retriever.
Another pre-retrieval strategy, query rewrite (Ma et al., 2023a), aims to close the gaps between the input text and the needed knowledge in retrieval, to reformulate the original question into a more conducive version to retrieve. Specifically, Ma et al. (2023a) propose the Rewrite-Retrieve-Read framework, which prompts an LLM to generate the query for the retrieval function. The motivation of the rewriting step is to clarify the retrieval need in the new query to ease the burden on the retrieval function to comprehend the input and enhance the output, i.e., retrieved relevant information. They have tested both the settings of using a frozen LLM and a trainable model to be the rewriter, both outperforming naive RAG or generation models, demonstrating diverse performance on different tested QA datasets though. Tan et al. (2024) also formulate a query rewriting strategy in their model that decomposes the heuristic answer from a proxy generation model into distinct claims.
@@
More recently, advanced RAG pipelines have been proposed using LLMs to generate reasoning paths and plans with the Information Retrieval (IR) module to iteratively retrieve knowledge to enhance LLM-based generation (Yao et al., 2023; Xu et al., 2023a; Shao et al., 2023). However, Zhu et al. (2023) point out that if the outputs of IR and LLM are low-quality, the retrieval and generation processes will get hindered by each other with such an iterative guidance pipeline. To overcome this barrier, they propose a new reasoning approach for query and retrieved knowledge enhancement. Post-retrieval strategies may also function to enhance the compatibility between the retrieved results and the generation models. For example, one of the main limitations of existing LLMs is the length of the input tokens, which prevents long retrieved documents being directly incorporated into existing RA-LLMs. For this limitation, Xu et al. (2023b) propose Retrieve, Compress, Prepend (RECOMP), which adds an intermediate step to process the retrieved documents into a textual summary before in-context augmentation in the generation process. From another perspective, long retrieved passage list leads to a high inference latency when using auto-regressive decoding at generation stage, which hurts the model’s efficiency. For this limitation, Hofstätter et al. (2023) propose a light version of FiD model that compresses the encoded vectors per retrieved passage before concatenating and feeding them through the decoder and also includes a re-ranker on the retrieved results before applying them in the generation.
Time Model Cite Retriever RetTrain RetAug Stage Pre-/Post-Retrieval Generator Aug Evaluation
2019 kNN-LM (Khandelwal et al., 2020) 619 DR(GP) No Inf RA DT Output LG
2020 REALM (Guu et al., 2020) 1437 DR(BE,BT) Yes PT+FT / ET Input OpenQA(NQ, WQ, CT)
2020 RAG (Lewis et al., 2020c) 2125 DR(DPR) Yes FT / ED (BART) Input OpenQA, AQA, Jeopardy QG, FV
2021 FiD (Izacard and Grave, 2021b) 780 SR(BM25)/ DR(DPR) No FT / ED (T5/BART) Input OpenQA
2021 SE-FiD (Komeili et al., 2022) 286 SE(Bing) No Inf RQG FiD Input WizInt, WoW
2021 FiD-KD (Izacard and Grave, 2021a) 190 DR(BE) Yes FT CR FiD Input OpenQA
2021 RETRO (Borgeaud et al., 2022) 683 DR(BERT, DPR) No PT / ED Inter LM, OpenQA
2021 EPR (Rubin et al., 2022) 384 DR(DPR) Yes Inf CR GPT-3,J,Neo, CODEX Demon UR
2022 OpenBook (Lazaridou et al., 2022) 145 SE+SR No QE GOPHER LM Input QA, FV
2022 DSP (Khattab et al., 2022) 117 ColBERTv2 No Inf RQG, RF GPT-3.5 Demon OpenQA, MHQA, CQA
2023 In-Context RALM (Ram et al., 2023) 211 DR/SR No Inf TRR GPT-2,J,Neo Input LM, OpenQA
2023 Atlas (Izacard et al., 2023) 367 DR(OE) Yes PT+FT / ED Input OpenQA, FV, WoW, EL,SF, MMLU
2023 FLARE (Jiang et al., 2023) 133 SR(BM25)/ SE(Bing) No Inf RQG GPT-3.5 Input MHQA, CR, LongQA, OS
2023 IRCoT (Trivedi et al., 2023) 114 SR(BM25) No Inf / GPT-3,Flan-T5 Input OpenQA
2023 Self-RAG (Asai et al., 2023b) 85 DR(OE) No FT CM tunable LLM OpenQA, LongQA, FV, BG
2023 REPLUG (Shi et al., 2023) 48 DR(BE) Yes FT TRA GPT-2,3 Inpput MMLU, OpenQA
2023 UDR (Li et al., 2023c) 42 DR(DPR) Yes FT CR GPT-Neo Demon 40 NLP tasks
2023 ITER-RETGEN (Shao et al., 2023) 40 DR(DPR) Yes FT RR InstructGPT, Llama-2 Input MHQA, FV, CR
Table 1. Basic information and main technical designs of high-impact RAG and RA-LLM models.1
11footnotetext: Retrievers: [BE: Bi-Encoder (also referred as dual encoder), OE: One-Encoder, BT: BERT-style Transformer, GP: Partial Generation, SE: Search Engine, SR: Sparse Retrieval, DPR:(Karpukhin et al., 2020)], Generators: [DT: Decoder-only Transformer, ET: Encoder-only Transformer, ED: Encoder-Decoder], Pre-/Post-Retrieval techniques: [RQG: Retrieval Query Generation, QE: Query Expansion, (T)RR: (Trainable) Re-Ranker, TRA: Trainable Retriever Adaptor, CR: Candidate Retrieval, CM: Critic Model], Augmentations:[Output: Output-layer Integration, Input: Input-layer Integration, Inter: Intermediate-layer Integration, Demon: As demonstration], Tasks: [AQA: Abstractive Question Answering, QG: Question Generation, NQ: Natural Questions, WQ: WebQuestions, CT: CuratedTrec, FV: Factor Verification, TQA: TriviaQA, WizInt: Wizard of the Internet task, WoW: Wizard of Wikipedia task, MHQA: Multi-hop QA, CQA: Conversational QA, EL: Entity Linking, SF: Slot-filling, MMLU: Massively-Multitask Language Understanding, CR: Commonsense Reasoning, LongQA: Long-form QA, OS: Open-domain Summarization, BG: Biography Generation, UR: Utterance Representing, RF: Retrieval Fusion]
@@ 3.1.4. Database
Retrieval in RAG is conducted based on external knowledge source, which can be a closed- or open-sourced (Ma et al., 2023a; Menick et al., 2022), as illustrated in Figure 3. Closed-sourced database generally stores key-value pairs for knowledge, which can be constructed in various ways. The keys are primarily used for similarity matching, being as sparse vectors such as in BM25 or dense embeddings from the retrieval encoding. The value depends on the specific retrieval target, which is raw text in most cases (Guu et al., 2020; Lewis et al., 2020c; Izacard and Grave, 2021b; Borgeaud et al., 2022; Lewis et al., 2020a; Seo et al., 2019). For example, each Wikipedia article is split into disjoint 100-word chunks, to make a total of 21M documents in early RAG (Lewis et al., 2020c). Each document is encoded by a dense embedding and saved in the database as the value and key, respectively. The value can store tokens too, one for each as applied in kNN-LM (Khandelwal et al., 2020) and SPALM (Yogatama et al., 2021). The source of the database depends on the specific application domains and tasks. Wikipedia is one of the most commonly applied general retrieval sets in previous RAG work, which stores factual structured information and has several versions differing in scale, from billion token-level (Khandelwal et al., 2020; Yogatama et al., 2021; Lewis et al., 2020c; Guu et al., 2020; Févry et al., 2020; de Jong et al., 2022; Xu et al., 2023b; Shi et al., 2023; Ram et al., 2023) to trillion token-level (Borgeaud et al., 2022). Domain-specific database is also used for downstream tasks. For example, for the code generation task, Zan et al. (2022) collect API information and code files of public libraries to build their APIretriever database. In addition, Zhou et al (Zhou et al., 2022) propose to use a documentation pool frequently updated with new content (newly released libraries) in their model.
@@
Applying Internet searching engine (Luo et al., 2023b) such as Bing and Google avoids the maintenance of the search index and can access up-to-date knowledge (Komeili et al., 2022; Lazaridou et al., 2022). Meanwhile, it provides a broader knowledge base than the closed-sourced database (Asai et al., 2023b; Lazaridou et al., 2022). It can also provide high-quality ranking after being tuned over decades of use. Internet search has been widely incorporated with black-box LLMs and shows effectiveness for different functions such as knowledge augmentation (Lazaridou et al., 2022), fact-checking (Menick et al., 2022) and LLM agent enhancement (Yao et al., 2023). Compared to traditional RAG, Internet search has been leveraged more as the retriever in RA-LLMs owing to the extraordinary capability of LLMs to be the Reader to comprehend the searching results, i.e., the retrieved documents, as well as LLMs’ ability to use tools to process and analyze the them (Ma et al., 2023a). Existing studies (Yu et al., 2023a) have shown that leveraging search engines (e.g., InstrucGPT) is particularly effective for LLMs on zero-shot knowledge-intensive tasks such as OpenQA and fact checking.
@@ 3.4. Retrieval Augmentation Necessity and Frequency
Most existing methods determine the necessity of retrieval based on the preliminary answers of LLMs or their internal reasoning results (Ram et al., 2023; Min et al., 2022). For example, Self-RAG (Asai et al., 2023b) introduces special tokens to assess the necessity of retrieval and control retrieval behavior. Other methods design iterative prompts to decide if extra information is required during generation, which thereby needs to invoke retrieval or other actions for LLMs (Yao et al., 2023; Wei et al., 2022). Wang et al. (2023a) propose Self-Knowledge guided Retrieval augmentation (SKR) method, which uses LLMs themselves or explicit small trainable models to offer self-knowledge as the reference for the adaptive calling of a retriever. In traditional RAGs, retrieval necessity judgment has also been explored and proposed to address by intuitive approaches such as assessing the confidence of the logits produced by the generation models (Jiang et al., 2021; Kadavath et al., 2022; He et al., 2021b). Such a solution is also applicable to RA-LLMs, for example, FLARE (Jiang et al., 2023) dynamically triggers RAG if logits are lower than a specific threshold. Tan et al. (2024) introduce a more flexible model SlimPLM, which detects missing knowledge in LLMs with a slim proxy model, which functions to generate a “heuristic answer”. The “heuristic answer” is used to assess the necessity of retrieval and facilitate the retrieval process for query rewriting when necessary.
@@ 4.1. Training-free
With the huge number of parameters, LLMs have exhibited human-level intelligence and achieved promising prediction performance on various downstream tasks. However, it is extremely challenging to frequently perform fine-tuning and update the knowledge stored in the model parameters (Lewis et al., 2020c) due to the considerable time and computational resources required. Recently, numerous studies have suggested enhancing LLMs with retrieval mechanisms, enabling them to dynamically acquire new knowledge from external sources without extra training processes (i.e., training-free(Izacard and Grave, 2021b; Jiang et al., 2023; Khattab et al., 2022), instead of relying solely on the implicit knowledge encoded in the model’s parameters. These approaches have shown significant performance improvement for various knowledge-intensive tasks, such as open-domain question answering (Lewis et al., 2020c). According to the different ways in which LLMs utilize retrieved information, we categorize these training-free methods into two categories: 1) Prompt Engineering-based Methods integrate retrieved knowledge into the original prompt directly, and 2) Retrieval-Guided Token Generation Methods retrieve information to calibrate the token generation process.
@@ 4.1.1. Prompt Engineering-based Methods
As the LLMs’ generation performance highly depends on the input query, numerous training-free RAG approaches employ external knowledge by refining the original prompts (Jiang et al., 2023; Khattab et al., 2022; Li et al., 2023d). Specifically, the retrieved texts are usually used as contextual information and combined with the original prompt to guide the generation of LLMs (Izacard and Grave, 2021b; Jiang et al., 2023; Khattab et al., 2022; Purwar and Sundar, 2023; Li et al., 2023d; Wang et al., 2023g; Kim et al., 2023). For example, In-Context RALM (Ram et al., 2023) keeps the LLM parameters unchanged and directly incorporates the retrieved document before the original prompt to augment the generation process. IRCoT (Trivedi et al., 2023) interleaves chain-of-thought (CoT) generation and knowledge retrieval steps, enabling the retrieval of more relevant information for subsequent reasoning steps compared to standard retrieval methods that rely solely on the question as the query. Instead of retrieving knowledge from a large corpus, GENREAD (Yu et al., 2023a) first prompts a LLM to generate contextual documents based on the query, and then generate answers based on the given context and question. SKR (Wang et al., 2023a) proposes guiding LLMs to determine whether they can answer a given question based on their internal knowledge, enabling flexible utilization of both internal and external knowledge by selectively calling the retriever. TOC (Kim et al., 2023) first retrieves relevant knowledge for ambiguous questions and recursively constructs a tree structure by clarifying ambiguous questions into multiple disambiguate questions, which is further aggregated to generate long-form answers.
@@ 5.1.1. QA Systems
QA Systems aim to provide precise answers to user’s queries. However, even when trained on extensive data, these systems may lack the latest information or specific domain knowledge that is not included in their training data (Izacard and Grave, 2021b; Liu et al., 2022c). To address this limitation, the integration of RA-LLMs has played a crucial role in advancing the capabilities of QA systems by enhancing their ability to retrieve and synthesize relevant information (Borgeaud et al., 2022; Izacard and Grave, 2021b). Specifically, RA-LLMs can provide coherent and contextually relevant answers by leveraging their retrieval component to access a vast knowledge base. For example, REALM (Guu et al., 2020) integrates a knowledge retriever that can retrieve information from a large corpus during pre-training, fine-tuning, and inference. This approach allows REALM to effectively retrieve from a vast knowledge corpus, thereby improving the accuracy of its responses. Similarly, Fusion-in-Decoder (Izacard and Grave, 2021b) retrieves passages from support documents and then fuses them with questions to generate the answer, achieving higher accuracy. In addition,  Borgeaud et al. (2022) indicate that the quality of the answers may rely more on the output of retrieval.
@@ 6. Future Challenges and Opportunities
Multi-Lingual RA-LLMs. The ability of leveraging knowledge from multiple languages can greatly enhance the capabilities of retrieval-augmented language models. As the world becomes increasingly interconnected, there is a growing need for AI systems that can understand and communicate across different languages. By incorporating multilingual knowledge retrieval and generation, these models can access and synthesize information from diverse linguistic sources, enabling more comprehensive and nuanced understanding and generation capabilities. Additionally, multilingual models can facilitate cross-cultural communication and knowledge sharing and breaking down language barriers, thereby bringing convenience to people across different regions of the world, especially those in areas with minority languages (Kabra et al., 2023; Li et al., 2023d). For example, users from countries with less prevalent languages can utilize abundant English and Chinese corpora for knowledge retrieval, enhancing the performance of large language models in downstream tasks.
@@
Quality of External Knowledge. As a commonly used datastore in current RAG systems, Wikipedia (Zhu et al., 2024; Karpukhin et al., 2020) serves as a vast repository of external textual knowledge used to augment the generation process, which contains millions of articles covering various disciplines. However, the reliability and accuracy of individual articles within Wikipedia vary significantly, and the introduction of some texts that deviate from facts might even mislead the model’s generation process. Therefore, it is crucial to enhance the quality of the external knowledge corpus and mitigate the negative impact of low-quality knowledge on the performance of LLMs. By enhancing the quality of the external knowledge and tailing robust mechanisms by filtering out low-quality or unreliable information, the RA-LLM systems might produce more accurate, reliable outputs, thereby improving their effectiveness in various real-world applications.